OpenStudy (anonymous):

a rectangular lawn has an area of 228 sq ft. you want to border it. the lawn is twice as long as it is wide. how much fencing do you need

OpenStudy (anonymous):

lets call one side \(x\) which makes the other side \(2x\) and the area is therefore \(x\times 2x=2x^2\) now we can solve because we know \(2x^2=228\)

OpenStudy (anonymous):

this tells us that \(x^2=144\) and so \(x=12\) perimeter is \(x+2x+x+2x=6x\) and since \(x=12\) we can find that number